1.Experience shows that the average number of misses by a darts player hitting off target is 5. What is the probabilty that on any darts night, a darts player will miss three times
2. Assume a sample of 25 elements taken from a population of 100 elements and the sample average was found to be 35.5 with a standard deviation of 10.Contruct 90% and 95% confidience intervals.

Answer by ewatrrr(24785)

1. P(x=3) = poissonpdf(5,3)
2. ME = 1.645(10/√25)
CI: = 35.5 - ME < mu < 35.5 + ME
ME = 1.96(10/√25)
CI: = 35.5 - ME < mu < 35.5 + ME